# A Persistence-Driven Edit Distance for Graphs with Abstract Weights

### v3

@@ Abstract
In this work we define a novel edit distance for graphs considered with some weights on the edges. The metric is driven by the idea of considering graphs as topological summaries in the context of persistence and topological data analysis. Several examples related to persistent sets are presented. In case the graphs are trees, the metric can be computed with a dynamical binary linear programming approach. This framework is applied and further studied in other works focused on merge trees, where the problems of stability and merge trees estimation are also assessed.
@@ 1 Introduction
The use of graphs is almost ubiquitous in science: any time some system needs to be discretised or some relationships between points/statistical units/objects formalized or represented, graphs are always involved.
For this reason a huge number of mathematical tools have been defined to work with graphs: differential calculus, distances between graphs, algebraic structures to solve graph-related problems and all other sorts of results have been developed throughout the years.
In this paper we consider graphs assuming the point of view of topological data analysis (TDA) (Edelsbrunner and Harer, 2008), i.e. we consider them as representing topological information, and develop an edit distance to analyse populations of graph-shaped topological summaries. TDA is a quite recent type of data analysis which resorts to algebraic topology (Hatcher, 2000) to extract from data information regarding each statistical unit’s shape. A typical TDA pipeline starts from a topological space and then induces a family of nested subspaces, called filtration, from which some topological information is extracted via the use of some functors, for instance simplicial homology functors (Hatcher, 2000). Generalizations of this standard procedure include using family of subspaces indexed by partially ordered sets (posets), like $\mathbb{R}^{n}$ ($n\geq 2$) (Botnan and Lesnick, 2022) and the use of Laplacians (Mémoli et al., 2022), path-connected components (Curry et al., 2022, 2021; Pegoraro, 2024a; Cavinato et al., 2022), zig-zag homology (Carlsson and De Silva, 2010), cosheaves (De Silva et al., 2016), families of functors (Turner et al., 2014) etc..
@@
Our investigation starts from objects of the form $F:(P,\leq)\rightarrow\operatorname*{Sets}$ (with $(P,\leq)$ being a poset), which generalize persistent sets as defined in Carlsson and Mémoli (2013), and arise from very general situations like $(P,\leq)\rightarrow\operatorname*{Top}\xrightarrow{\pi_{0}}\operatorname*{Sets}$ (with $\operatorname*{Top}$ being the category of topological spaces and $\pi_{0}$ being the functor of path-connected components). For instance, if $(P,\leq)$ is given by the ordered set $\{f^{-1}((-\infty,t])\}_{t\in\mathbb{R}}$ for some $f:X\rightarrow\mathbb{R}$ we recover merge trees of real valued functions (Morozov et al., 2013; Curry et al., 2021; Pegoraro, 2024a), while if $(P,\leq)$ is the poset of open intervals in $\mathbb{R}$ we recover Reeb cosheaves (De Silva et al., 2016).
@@
Relationships between posets and graphs have been extensively studied: for instance it is well-known that any poset can be represented with a directed acyclic graph (DAG) (Aho et al., 1972). Thus, it is no surprise that a common way to represent instances of $F:(P,\leq)\rightarrow\operatorname*{Sets}$ is via some kind of graphs, like merge trees or Reeb graphs.
Metrics to compare graphs are not a novel research topic (Billera et al., 2001; Gao et al., 2010; Bille, 2005; Feragen and Nye, 2020), but recent advances in stratified spaces, optimal transport and geometric data analysis have fostered some novel techniques to deal with such objects (Mémoli, 2011; Calissano et al., 2023). Similarly, the use of graphs (possibly labeled Yan et al. (2019)) as summaries of persistent sets has produced a series of original definitions and results (Bauer et al., 2014; De Silva et al., 2016; Di Fabio and Landi, 2016; Bauer et al., 2020; Morozov et al., 2013; Gasparovic et al., 2019; Pont et al., 2022; Sridharamurthy et al., 2020; Wetzels et al., 2022; Sridharamurthy and Natarajan, 2021; Cardona et al., 2021; Touli, 2020; Curry et al., 2022) which are all driven by these novel scenarios.
To differentiate this work from the previous ones, we quickly mention that the metric here presented, when used (up to some minor modifications) to compare merge trees, satisfies some stability properties (contrary to Pont et al. (2022); Sridharamurthy et al. (2020); Wetzels et al. (2022); Sridharamurthy and Natarajan (2021)), which are analogous to the ones of the $1$-Wasserstein distance between persistence diagrams (Edelsbrunner and Harer, 2022). In particular our metric adds all the local modifications needed to turn one graph into the other, while the other stable metrics look for the biggest modification that needs to be made, making those metrics more stable, but also less sensitive to differences on a global scale. Exactly as for the bottleneck distance and the Wasserstein distances between persistence diagrams. Formal details are given in Pegoraro (2024a). Lastly, differently from the other stable metrics, we provide, implement and employ an exact algorithm to compute the metric for trees.
@@
In this paper we define a novel edit distance apt to compare graphs which are interpreted as topological summaries. Since, increasingly often, such objects are enriched with some kind of additional information which is, to some extent, attached to vertices and edges of the graphs (Ushizima et al., 2012; Curry et al., 2022; Pegoraro, 2024b; Curry et al., 2023a), we keep our dissertation as general as possible, allowing for some abstract weights to be associated to the edges in the graph. In particular, we can articulate the contributions of this paper in two distinct points:
@@
    the main contributions are given by the definition of a distance between weighted graphs, with weights lying in certain metric spaces, and a computational framework to exactly compute the distance with binary linear optimization when the graph is a tree;
@@
    we start exploring the possibilities offered by the different weights we can attach to the edges, discussing the potential applications for these abstract weights and how can they be used to tackle different data analysis scenarios (some of which have already been investigated in other works).
The edit distance we define stems from classical edit distances (Levenshtein et al., 1966; Tai, 1979; Bille, 2005; Gao et al., 2010) but add some deep modifications, which can be intuitively motivated by the following example, represented in Figure 1.
@@
Consider the following function: $f(x)=x\cdot\sin(x)$ for $x\in[0,2\pi]$ Figure 1(a). Then take the sublevel set filtration $F(t)=f^{-1}((-\infty,t])$ and the persistent set $\pi_{0}\circ F$. The information of this persistent set can be represented with a tree with four vertices and three edges, as in Figure 1(a), along with a real-valued height function defined on its vertices - with the root being at height +$\infty$+ ∞. We can associate to any edge its length, obtained as the difference in height between its endpoints. Suppose that we were to remove from the (merge) tree the edge $(a,b)$, we would obtain the red (merge) tree in Figure 1(c). Vertex $a$ is now an order $2$ vertex which represents a merging between path connected components (i.e. a local maximum in the function), which is no more happening. On a function level, this can be visualized as removing the local minimum in $2\pi$, for instance restricting the function on $[0,3/2\pi]$ Figure 1(b). Usual edit distances would regard the red and black weighted trees in the green box in Figure 1(c) as very different trees, having a different number of edges and very different weights associated to the edges. As a consequence also the black trees in the green boxes in Figure 1(a) and in Figure 1(b) would be very far apart. In our case however, we would like our metric to be much less sensitive and, for instance, not to distinguish between the black and red trees in trees Figure 1(c). For this reason, differently from other edit distances, we add the possibility to remove the vertex $a$ and join the two adjacent edges.
@@
As already pointed out, the framework we introduce here lies the foundations for a series of developments which are carried out in separate works. Some of these developments, in turn, contain results which retrospectively validate the approach we pursue in this manuscript. In particular, we have the following group of interconnected works:
  • the present work, where the main theoretic and computational framework is established, with the aim of providing a tool to work with graphs with general weights;
  • when the weights on the edges are just positive numbers, this general framework is used by Pegoraro (2024a) to define a metric for merge trees, which can be computed with the algorithm we present here. With the help of some results contained in Pegoraro and Secchi (2024), which is more focused on the statistical aspects of such metric, Pegoraro (2024a) shows that the approach we pursue here induces a metric with stability properties which are analogous to the ones of the $1$-Wasserstein distance between persistence diagrams. In Pegoraro and Secchi (2024) the metric is also tested (using the algorithm provided in this work) on a classification pipeline for a data set which is benchmark in the field of functional data analysis; ;
  • Pegoraro (2024b), instead, exploits the present work to explore the case with weights being functions defined on the edges of the trees. The stability result obtained in Pegoraro (2024b) stems from the one proved in Pegoraro and Secchi (2024);
  • finally, Cavinato et al. (2022) considers hierarchical clustering dendrograms and carries out an analysis of a radiomic data set, putting to work the framework (theoretic and computational) we define here on real data. Therein, a stability result in terms of Gromov-Hausdorff distance between point clouds is also obtained (which relies again on the one in Pegoraro and Secchi (2024)).
@@
Section 2 introduces the idea of editable spaces, which are the spaces we want the weights of our graphs to live in. Their properties are exploited in Section 3 to define editable graphs. Section 4 develops the novel edit distance and proves that it is well defined for graphs considered up to order $2$ vertices. Section 5 presents a series of general examples connecting persistent sets, and their generalizations, to editable graphs. Section 6 specialises the previously introduced tools to the case where the graph is a tree and obtains an algorithm to compute the distance via linear binary optimization. Section 7 concludes the paper by highlighting pros and cons of the approach along with possible further developments.
@@ 2 Editable Spaces
As already stated, the aim of the manuscript is to present a novel metric structure to work with weighted (and possibly directed) graphs, where the word weighted is intended in a broad sense: weights are given by a map associating the edges of a graph to elements in a metric space.
In order to define the metric, we need these spaces of weights to satisfy some properties which are fundamental to ensure some desired behaviours when modifying graphs.
@@
    $(X,\odot,0_{X})$ is a monoid (that is $X$ has an associative operation $\odot$ with zero element $0_{X}$)
@@
Exploiting the monoidal structure of an editable space we can define a partial order which we will use in later sections.
@@ 2.1.2 $1$-Wasserstein Space of Persistence Diagrams
Similarly $D(\Delta)$ is the set of formal sums of elements in $\Delta=\{(b,b)\in R^{2}\mid 0\leq b\}$. In both sets of formal sums we have an operation given by the sum of formal sums, which is commutative and has the empty formal sum as zero element. Given $p,q\in D(\mathbb{R}^{2}_{\Delta})$ we say that $s\sim q$ if there exist $r,t\in D(\Delta)$ such that $p+r=q+t$. We collect the equivalence classes of such relationships in $D(\mathbb{R}^{2}_{\Delta},\Delta)$ and call them persistence diagrams. The sum operation and the zero element are coherently defined on the quotient.
As a consequence of 2 we can write every diagram $D$ as a sum of points in $\mathbb{R}^{2}_{\Delta}$ each with coefficient $1$. This “decomposition” is clearly non-unique as it varies between elements in the same equivalence class. With an abuse of notation we may say that $x\in D$, meaning that $x=(b,d)$ is an off-diagonal point of $\mathbb{R}^{2}_{\Delta}$ (i.e. $d>b$) appearing in such decomposition of an element (in fact, in any element) in the equivalence class of $D$.
@@
$W_{1}(D,D^{\prime})=\inf_{\gamma:D\rightarrow D^{\prime}}\parallel x-\gamma(x)\parallel_{1},$
where $\gamma$ is any bijection between an element in the equivalence class of $D$ and an element in the equivalence class $D^{\prime}$. The infimum in the definition of $W_{1}(D,D^{\prime})$ is always attained by some $\gamma$, and thus, it is always a minimum.

### v4

# A Persistence-Driven Edit Distance for Trees with Abstract Weights

@@ Abstract
In this work we define a novel edit distance for trees considered with some abstract weights on the edges. The metric is driven by the idea of considering trees as topological summaries in the context of persistence and topological data analysis. Several examples related to persistent sets are presented. The metric can be computed with a dynamical binary linear programming approach. This framework is applied and further studied in other works focused on merge trees, where the problems of stability and merge trees estimation are also assessed.
@@ 1 Introduction
Graphs are widely used across scientific disciplines. Whenever a system is discretized, or when relations among points, statistical units, or abstract objects need to be formalized, graphs naturally emerge as a fundamental modeling tool.
This ubiquity has led to the development of a broad array of mathematical techniques for analyzing and comparing graphs, ranging from differential calculus on discrete structures and graph metrics, to specialized algebraic frameworks. Over time, this has given rise to a rich and diverse landscape of graph-theoretic methods.
Within this broader context, trees occupy a particularly important role. Their structural simplicity, combined with considerable expressive power, offers an appealing balance between interpretability and computational tractability. As a result, trees have become central in many areas of science. Two especially prominent examples are phylogenetic trees Billera et al. (2001); Felsenstein and Felenstein (2004); Garba et al. (2021); Lueg et al. (2024) and clustering dendrograms Murtagh and Contreras (2017); Xu and Tian (2015), which are closely related structures. Both describe how a set of elements evolves or merges under some notion of similarity or agglomeration, and both illustrate the versatility of trees as a means of encoding hierarchical information.
In this paper, we approach trees from the perspective of topological data analysis (TDA) (Edelsbrunner and Harer, 2008), interpreting them as carriers of topological information. Our goal is to develop an edit distance tailored to analyzing populations of tree-shaped topological summaries.
TDA is a relatively recent framework for data analysis that leverages tools from algebraic topology (Hatcher, 2000) to extract shape-related features from data at the level of individual statistical units.
A typical TDA pipeline begins with a topological space and constructs a nested sequence of subspaces—known as a filtration—from which topological information is extracted using functors such as simplicial homology (Hatcher, 2000). This standard approach has been generalized in several directions, including filtrations indexed by more general partially ordered sets (e.g., $\mathbb{R}^{n}$ for $n\geq 2$) (Botnan and Lesnick, 2022), and the use of other mathematical objects such as Laplacians (Mémoli et al., 2022), path-connected components (Curry et al., 2022, 2021; Pegoraro, 2024a; Cavinato et al., 2022), zigzag homology (Carlsson and De Silva, 2010), cosheaves (De Silva et al., 2016) among others.
Previous Works on Edit Distances for Trees in TDA
@@
The topological perspective we assume in this work stems from persistent sets, as defined in Carlsson and Mémoli (2013), which are functors of the form $F:(P,\leq)\rightarrow\operatorname*{Top}\xrightarrow{\pi_{0}}\operatorname*{Sets}$ (with $(P,\leq)$ being a poset, $\operatorname*{Top}$ being the category of topological spaces and $\pi_{0}$ being the functor of path-connected components). The most typical example being $(P,\leq)$ equal to $\{f^{-1}((-\infty,t])\}_{t\in\mathbb{R}}$ for some $f:X\rightarrow\mathbb{R}$, which amounts to using merge trees of real valued functions (Morozov et al., 2013; Curry et al., 2021), while, if $(P,\leq)$ is the poset of open intervals in $\mathbb{R}$, up to some gluing conditions, we recover Reeb cosheaves (De Silva et al., 2016).
@@
Relations between posets and graphs have been extensively studied: for instance it is well-known that any poset can be represented with a directed acyclic graph (DAG) (Aho et al., 1972). Thus, it is no surprise that a common way to represent instances of $F:(P,\leq)\rightarrow\operatorname*{Sets}$ is via some kind of graphs, like merge trees or Reeb graphs. Accordingly, the use of graphs (possibly labeled Yan et al. (2019)) as summaries of persistent sets has produced a series of original definitions and results (Bauer et al., 2014; De Silva et al., 2016; Di Fabio and Landi, 2016; Bauer et al., 2020; Morozov et al., 2013; Gasparovic et al., 2019; Pont et al., 2022; Sridharamurthy et al., 2020; Wetzels et al., 2022; Sridharamurthy and Natarajan, 2021; Cardona et al., 2021; Touli, 2020; Curry et al., 2022) which are all driven by these novel scenarios.
A significant number of these works focus on defining suitable metric structures for merge trees and Reeb graphs through variants of edit distances. In particular, (Di Fabio and Landi, 2016; Bauer et al., 2020) are focused on sup-norm kinds of distances, where the difference between (Reeb) graphs is determined by the cost of the biggest modification needed to turn one graph into another. While Pont et al. (2022); Sridharamurthy et al. (2020); Wetzels et al. (2022); Sridharamurthy and Natarajan (2021) consider edit distances for merge trees which are more closely related to classical tree edit distances Tai (1979), where the objective is to minimize the total cost of a sequence of modifications converting one merge tree into another.
In this work, we propose an edit distance for weighted trees that aligns more closely with the second family of approaches mentioned above, but it differs from previous methods in two key ways:
  • Stability Properties: in Pegoraro (2024a) and Pegoraro and Secchi (2024) it is shown that, when adapted to merge trees, the metric we define here produces an edit distance $d_{E}$ satisfying the following stability relation :
    $d_{I}(T,T^{\prime})\leq d_{E}(T,T^{\prime})\leq 2(\text{size}(T)+\text{size}(T^{\prime}))d_{I}(T,T^{\prime})$
    where $d_{I}$ is interleaving distance between merge trees (Morozov et al., 2013) and $\text{size}(T)$ is the number of edges in $T$. Which is analogous to the relation between the bottleneck distance ($d_{B}$) and the $1$-Wasserstein metric ($W_{1}$) between persistence diagrams (PDs):
    $d_{B}(D,D^{\prime})\leq W_{1}(D,D^{\prime})\leq(\#D+\#D^{\prime})d_{B}(D,D^{\prime}).$
    See Edelsbrunner and Harer (2010). Among the edit distances presented in Pont et al. (2022); Sridharamurthy et al. (2020); Wetzels et al. (2022); Sridharamurthy and Natarajan (2021), only the one in Wetzels et al. (2022) was shown—in Pegoraro (2024a)—to posses some stability properties. This stability result also clarifies the theoretical distinction between our approach and those in Bauer et al. (2014); Di Fabio and Landi (2016); Bauer et al. (2020), all of which yield distances that are bounded from above by $d_{I}$. Since $d_{B}$ is the universal metric between PDs and $d_{I}$ is the universal one between merge trees, by analogy, one can further characterize the difference between $d_{E}$ and Di Fabio and Landi (2016); Bauer et al. (2020), via the easier and well understood difference between $d_{B}$ and $W_{1}$.
  • Abstract Weights: we investigate generalizations in which trees are enriched not only with edge lengths, but with weights taking values in more general metric spaces. This allows topological summaries to capture and represent a wider variety of information.
@@
As anticipated, the main contribution of this work is the introduction of a novel edit distance designed to compare trees interpreted as topological summaries. Increasingly, such trees are enriched with additional information, typically associated with the vertices or edges of the graph, as seen in various recent works (Ushizima et al., 2012; Curry et al., 2022; Pegoraro, 2024b; Curry et al., 2023a). To remain as general as possible, we allow abstract weights to be assigned to the edges of the trees.
Given the wide range of settings that emerge from this enriched perspective (as discussed in Section 5 and reflected in the diverse works built on the framework we establish) this paper focuses on the following core contributions:
@@
    Defining a general edit distance for trees with abstract weights;
@@
    Formalizing the abstract properties that edge weights must satisfy for the metric to be well-defined;
  3. 3.
    Developing a computational framework for efficiently computing the metric;
  4. 4.
    Beginning to explore the variety of edge weights that can be used, discussing their potential applications and how can they support diverse data analysis scenarios. We leave a deeper investigation into the interpretability of the metric in specific scenarios and the study of its stability properties to other works.
The edit distance we define stems from classical edit distances (Levenshtein et al., 1966; Tai, 1979; Bille, 2005; Gao et al., 2010; Lerouge et al., 2016), but add some deep modifications, which can be intuitively motivated by the following example, represented in Figure 1.
@@
Consider the following function: $f(x)=x\cdot\sin(x)$ for $x\in[0,2\pi]$ Figure 1(a). Then take the sublevel set filtration $F(t)=f^{-1}((-\infty,t])$ and the persistent set $\pi_{0}\circ F$. The information of this persistent set can be represented with a tree with four vertices and three edges, as in Figure 1(a), along with a real-valued height function defined on its vertices - with the root being at height +$\infty$+ ∞. We can associate to any edge its length, obtained as the difference in height between its endpoints. Suppose that we were to remove from the (merge) tree the edge $(a,b)$, we would obtain the red (merge) tree in Figure 1(c). Vertex $a$ is now a degree $2$ vertex which represents a merging between path connected components (i.e. a local maximum in the function), which is no more happening. On a function level, this can be visualized as removing the local minimum in $2\pi$, for instance restricting the function on $[0,3/2\pi]$ Figure 1(b). Usual edit distances would regard the red and black weighted trees in the green box in Figure 1(c) as very different trees, having a different number of edges and very different weights associated to the edges. As a consequence also the black trees in the green boxes in Figure 1(a) and in Figure 1(b) would be very far apart. In our case however, we would like our metric to be much less sensitive and, for instance, not to distinguish between the black and red trees in trees Figure 1(c). For this reason, differently from other edit distances, we add the possibility to remove the vertex $a$ and join the two adjacent edges.
@@
The framework introduced in this work lays the foundation for a series of developments explored in separate studies: the edit distance we define is applied across different contexts, each accompanied by dedicated theoretical results that both justify the specific application and reinforce the foundational framework developed in this manuscript.
We summarize such relations as follows:
  • All the works we mention in the following points rely heavily on Section 4.2 and 2 for developing theoretic results, and on Algorithm 1 for applications and case studies.
  • Merge trees: this general framework is used by Pegoraro (2024a) to define a metric for merge trees, which can be computed with Algorithm 1. With the help of some results contained in Pegoraro and Secchi (2024), which is more focused on the statistical aspects of such metric, Pegoraro (2024a) shows that the approach we pursue here induces a metric with stability properties which are analogous to the ones of the $1$-Wasserstein distance between persistence diagrams. Pegoraro and Secchi (2024) and Cavinato et al. (2022) present real data applications to show the practical effectiveness of the metric.
  • Functions defined on merge trees: Pegoraro (2024b) exploits the present work, exploring the case where weights are functions defined on the edges of the trees, obtaining another stability result.
@@
Section 2 introduces the idea of editable spaces, which are the spaces we want the weights of our graphs to live in. Their properties are exploited in Section 3 to define dendrograms. Section 4 develops the novel edit distance and proves that it is well defined for trees considered up to degree $2$ vertices. Section 5 presents a series of general examples regarding persistent sets. Section 6 obtains an algorithm to compute the distance via linear binary optimization. Section 7 concludes the paper by highlighting the pros and cons of the approach along with possible further developments, with a brief comparison with graph edit distances.
@@ 2 Editable Spaces
As outlined in the introduction, the goal of this manuscript is to introduce a novel metric structure for working with weighted trees, where the term weighted is interpreted broadly: edge weights are defined via a map that assigns each edge a value in a metric space. To define this metric meaningfully, we require the metric spaces to satisfy certain properties. These conditions are essential to guarantee well-behaved transformations when modifying trees.
@@
    $(X,\odot,0_{X})$ is a commutative monoid (that is $X$ has an associative operation $\odot$ with zero element $0_{X}$)
@@
Exploiting the monoidal structure of an editable space, we can turn it into a partially ordered set.
@@
If the editable space $X$ is contained in a group $V$, the partial order can be further characterized.
@@ 2.1.2 $1$-Wasserstein Space of Persistence Diagrams
Similarly $D(\Delta)$ is the set of formal sums of elements in $\Delta=\{(b,b)\in\mathbb{R}^{2}\mid 0\leq b\}$. In both sets of formal sums we have an operation given by the sum of formal sums, which is commutative and has the empty formal sum as zero element. Given $D_{1},D_{2}\in D(\mathbb{R}^{2}_{\Delta})$ we say that $D_{1}\sim D_{2}$ if there exist $D_{3},D_{4}\in D(\Delta)$ such that $D_{1}+D_{3}=D_{2}+D_{4}$. We collect the equivalence classes of such relations in $D(\mathbb{R}^{2}_{\Delta},\Delta)$ and call them persistence diagrams, and indicate such an equivalence class with the notation $[D]$. The sum operation and the zero element are coherently defined on the quotient.
As a consequence of 2 we can write every diagram $[D]$ as a sum of points in $\mathbb{R}^{2}_{\Delta}$ each with coefficient $1$. This “decomposition” is clearly non-unique as it varies between elements in the same equivalence class. With an abuse of notation we may say that $x\in[D]$, meaning that $x=(b,d)$ is an off-diagonal point of $\mathbb{R}^{2}_{\Delta}$ (i.e. $d>b$) appearing in such decomposition of an element (in fact, in any element) in the equivalence class of $D$.
@@
$W_{1}([D],[D^{\prime}])=\inf_{\gamma:\widetilde{D}\rightarrow\widetilde{D}^{\prime}}\parallel x-\gamma(x)\parallel_{1},$
where $\gamma$ is any bijection between some $\widetilde{D}$ such that $\widetilde{D}\sim D$, and some $\widetilde{D}^{\prime}$ such that $\widetilde{D}^{\prime}\sim D^{\prime}$. The infimum in the definition of $W_{1}([D],[D^{\prime}])$ is always attained by some $\gamma$, and thus, it is always a minimum.
@@
Remark 1
We assumed 3 as it makes 2 easier to prove, compared to other formulations. However, for the sake of simplicity, we drop the notation $[D]$ in the remaining of the manuscript referring to a diagram simply as $D$, chosing the representative with no points of the form $(b,b)$.